Describe the translations applied to the graph of y 5 x2 to obtain a K graph of each quadratic function. Sketch the graph.

y=3x^2-4
1
Expert's answer
2020-12-01T09:32:21-0500


But, to make sure you're up to speed, a parabola is a type of U-Shaped curve that is formed from equations that include the term x 2 x^{2} x2. Oftentimes, the general formula of a quadratic equation is written as: y = ( x − h ) 2 + k y = (x-h)^{2} + k y=(x−h)2+k.